OpenStudy (anonymous):

Simplify radical 10 times radical 8

jimthompson5910 (jim_thompson5910):

How can you factor 8 so it has one factor that's a perfect square?

OpenStudy (anonymous):

would radical 8 turn into 2 radical 2?

jimthompson5910 (jim_thompson5910):

good, do you see how \[\Large \sqrt{8} = 2\sqrt{2}\]

OpenStudy (anonymous):

yes, so when i try to simplify the two together how is that done?

jimthompson5910 (jim_thompson5910):

you would use the rule \[\Large \sqrt{x}*\sqrt{y} = \sqrt{x*y}\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{10}*\sqrt{8}\] \[\Large \sqrt{10}*2*\sqrt{2}\] \[\Large 2*\sqrt{10}*\sqrt{2}\] \[\Large 2*\sqrt{10*2} \ ... \ \text{Use the rule } \sqrt{x}*\sqrt{y} = \sqrt{x*y}\] \[\Large 2*\sqrt{20}\] What's next?

OpenStudy (anonymous):

its it become 2 radical 5?

jimthompson5910 (jim_thompson5910):

correct, \[\Large \sqrt{20} = 2\sqrt{5}\] making the final answer to be what?

OpenStudy (anonymous):

2 radical 5 is not the final answer?

jimthompson5910 (jim_thompson5910):

no, it helps you get there though

jimthompson5910 (jim_thompson5910):

do you see how \[\Large \sqrt{20} = 2\sqrt{5}\] ??

OpenStudy (anonymous):

yes, i see how we get that but how does this get to the final answer? Am i forgetting a number somewhere?

jimthompson5910 (jim_thompson5910):

if you look back up at my steps, you'll see the last step is \[\Large 2*\sqrt{20}\] so you're forgetting that 2* out front

OpenStudy (anonymous):

oh, sorry so its is 4 radical 5?

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

an alternative way is to multiply the radicands to get 10*8 = 80, then simplify \[\large \sqrt{80}\]

OpenStudy (anonymous):

Thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

I prefer \[ \sqrt{10}\sqrt{8}=\sqrt{80}=\sqrt{(16) 5}=4\sqrt 5 \]

OpenStudy (anonymous):

Thank you for showing another way @eliassaab

OpenStudy (anonymous):

YW @Marie18